# Totally elliptic surface group representations

### v1

@@ 1.4. Acknowledgments
2. Proof of Theorem A
@@ 2.1. Overview
The proof of Theorem A unfolds over the next subsections. We start by showing that when the genus of $\Sigma$ is at least $1$, then the only totally elliptic representations are orthogonal representations (Section 2.2). The arguments mostly rely on the observation that a commutator of elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is itself elliptic only when it’s trivial (Lemma 2.1). We then briefly recall what DT representations are (Section 2.3) and prove that a totally elliptic genus-0 surface group representation that isn’t orthogonal must be a DT representation (Sections 2.3.12.3.3).
@@ 2.2. Positive genus
We start by proving that a totally elliptic representation $\rho\colon\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ of a surface $\Sigma$ of genus $g\geq 1$ is necessarily an orthogonal representation (Propositions 2.5 and 2.8). Most of the reasoning will rely on the following standard results about $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, which we include for the sake of completeness.
According to Definition 1.1, the elliptic elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ fall into two categories: the regular elliptic elements and the identity. The regular elliptic elements are those whose trace is smaller than $2$ in absolute value. Geometrically, they act on the hyperbolic plane $\operatorname{\mathbb{H}}$ by rotations around their unique fixed point. The subset of regular elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is homeomorphic to the open ball $(0,2\pi)\times\operatorname{\mathbb{H}}$, where the first factor records the rotation angle of a regular elliptic and the second its unique fixed point (the centre of rotation).
Lemma 2.1.
A regular elliptic element $A$ of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is contained in a unique maximal compact subgroup which consists of all the elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ which fix the unique fixed point of $A$. Furthermore, if $X$ is any element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, then the commutator
@@
is either the identity or a hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. The commutator is trivial if and only if $X$ is elliptic and fixes the unique fixed point of $A$.
@@
The unique maximal compact subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ that contains $A$ is its centralizer, which is conjugate to $\operatorname{PSO}(2)$. Since $A$ is regular, any element that commutes with $A$ is either the identity or a regular elliptic with the same fixed point as $A$.
Let’s now consider the commutator $[A,X]$. As we just said, if $[A,X]=1$, then $X$ is elliptic and fixes the unique fixed point of $A$. If $[A,X]\neq 1$ and $o$ denotes the unique fixed point of $A$, then $XA^{-1}X^{-1}$ is regular elliptic with fixed point $X.o\neq o$. Note that if $\alpha$ is the rotation angle of $A$, then the rotation angle of $XA^{-1}X^{-1}$ is $2\pi-\alpha$. Let $\ell$ denote the unique geodesic line in $\operatorname{\mathbb{H}}$ through $o$ and $X.o$. Similarly, let $\ell_{1}$ and $\ell_{2}$ denote the geodesic lines obtained by rotating $\ell$ anti-clockwise by an angle $\alpha/2$ around $o$, respectively $X.o$.
@@
Lemma 2.1 plays a central role in our discussion because it tells us that the commutator of two elliptic elements is elliptic if and only if it is trivial. It turns out that the commutator of two simple closed curves (seen as elements of $\pi_{1}\Sigma$) sometimes represents a simple closed curve too. We must, however, be careful, because a simple closed curve on $\Sigma$ is represented by many elements in $\pi_{1}\Sigma$, all conjugate to each other.
@@
If two simple closed curves on $\Sigma$ intersect once (their geometric intersection number is 1) and we represent them in the simplest way by two fundamental group elements $a$ and $b$ with basepoint being the intersection point, then $[a,b]$ also represents a simple closed curve on $\Sigma$.
@@
Pick two loops on $\Sigma$ that only intersect at the basepoint and whose homotopy classes are $a$ and $b$. Enlarge each loop by replacing it with a thin ribbon. The boundary of the union of the two ribbons is an embedded circle in $\Sigma$ which is in the free homotopy class determined by the fundamental group element $[a,b]$. ∎
As we’re about to see, Lemmas 2.1 and 2.2 together are the main obstruction for constructing totally elliptic surface group representations. Let’s start by introducing some notation. We pick a geometric generating family for $\pi_{1}\Sigma$ which consists in $2g+n$ generators $a_{1},\ldots,a_{g}$, $b_{1},\ldots,b_{g}$, and $c_{1},\ldots,c_{n}$ satisfying the unique relation
@@
A possible choice of geometric generators are the homotopy classes of the loops illustrated on Figure 1.
@@
We adopt the convention of concatenating loops by starting with the rightmost loop. More precisely, if $\gamma_{1}$ and $\gamma_{2}$ are two loops on $\Sigma$ based at the same point, then their concatenation $\gamma_{1}\gamma_{2}$ is the loop obtained by going along $\gamma_{2}$ first, and then along $\gamma_{1}$. This rule applies to fundamental group multiplications too.
@@
If the genus $g$ of $\Sigma$ is at least 2 and $i\neq j$ are two indices in $\{1,\ldots,g\}$, then the following fundamental group elements represent simple closed curves on $\Sigma$:
@@
Each pair of fundamental group elements appearing in a given commutator represents two simple closed curves on $\Sigma$ with geometric intersection number $1$. So, by Lemma 2.2, we can expect their commutators to represent simple closed curves too. We must, however, be careful when we represent a simple closed curve by a fundamental group element because of basepoint issues. The choice we made ensures that each commutator indeed represents a simple closed curve on $\Sigma$. This can be seen by drawing the corresponding curves on Figure 1. ∎
@@
Assume that $\Sigma$ is closed and has genus $g\geq 1$. If $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is a totally elliptic representation, then $\rho$ is an orthogonal representation.
@@
We’ll write $A_{i}=\rho(a_{i})$ and $B_{i}=\rho(b_{i})$. If $g=1$, then the image of $\rho$ is generated by $A_{1}$ and $B_{1}$ which satisfy $[A_{1},B_{1}]=1$. So, the image of $\rho$ is generated by two commuting elliptic elements, hence contained in a compact subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
Let’s now consider the case $g\geq 2$. If $\rho$ is the trivial representation, then it’s orthogonal. If not, then there exists an index $i\in\{1,\ldots,g\}$ such that either $A_{i}\neq 1$ or $B_{i}\neq 1$. We’ll look at the case where $A_{i}\neq 1$ (the case where $B_{i}\neq 1$ is identical). Let $j\neq i$ be another index. Because $\rho$ is totally elliptic, it maps all the simple closed curves from Lemma 2.4 to the identity because of Lemma 2.1. In particular, we have $[A_{i},B_{i}]=1$ and $[A_{i},B_{i}^{-1}B_{j}^{-1}]=1$, implying that $[A_{i},B_{j}]=1$. Since $[A_{i},B_{i}]=1$, up to conjugating $\rho$, we may assume that $A_{i}$ and $B_{i}$ belong to $\operatorname{PSO}(2)$. From $[A_{i},B_{j}]=1$ and our assumption that $A_{i}\neq 1$, we conclude that $B_{j}\in\operatorname{PSO}(2)$ by Lemma 2.1.
We also infer from Lemmas 2.1 and 2.4 that $[A_{j},B_{j}]=[A_{j},B_{j}^{-1}B_{i}^{-1}]=[A_{j},B_{i}A_{i}^{-1}B_{j}]=1$. If $B_{j}\neq 1$, then we immediately conclude that $A_{j}\in\operatorname{PSO}(2)$ from $[A_{j},B_{j}]=1$ by Lemma 2.1. Since $[A_{j},B_{j}]=1$ and $[A_{j},B_{j}^{-1}B_{i}^{-1}]=1$, we deduce that $[A_{j},B_{i}]=1$ which also implies $A_{j}\in\operatorname{PSO}(2)$ as soon as $B_{i}\neq 1$. If both $B_{i}=1$ and $B_{j}=1$, then we alternatively observe that $[A_{i}^{-1},A_{j}]=[B_{i}A_{i}^{-1}B_{j},A_{j}]=1$ which gives $A_{j}\in\operatorname{PSO}(2)$ as well, because $A_{i}\neq 1$.
@@
The first fundamental group element in (2.1) can be rewritten as $b_{i}a_{i}^{-1}b_{i}^{-1}\cdot c_{j}^{-1}a_{i}c_{j}$, which is the product of a conjugate of $a_{i}^{-1}$ and a conjugate of $a_{i}$. A similar observation holds for the second fundamental group element in (2.1) which is equal to $b_{i}a_{i}b_{i}a_{i}^{-1}b_{i}^{-1}\cdot c_{j}^{-1}b_{i}^{-1}c_{j}$ and is thus a product of conjugates of $b_{i}$ and $b_{i}^{-1}$. By drawing a rapid picture, we convince ourselves that both represent simple closed curves on $\Sigma$. We prove in the same way that the fundamental group element (2.2) represents a simple closed curve too. ∎
@@
A surface group representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is said to be reduced if it doesn’t map any peripheral curve on $\Sigma$ to the identity. In terms of the geometric generators from Figure 1, a representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is reduced if it doesn’t map any generator $c_{i}$ to the identity.
Note that every non-reduced totally elliptic surface group representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ corresponds to a reduced totally elliptic representation $\pi_{1}\Sigma^{\prime}\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, where $\Sigma^{\prime}$ is obtained from $\Sigma$ by removing the punctures whose peripheral curves are mapped to the identity.
@@
Assume that $\Sigma$ is a surface of genus $g\geq 1$ and with $n\geq 1$ punctures. If $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is a reduced totally elliptic representation, then $\rho$ is an orthogonal representation.
@@
As before, we’ll write $A_{i}=\rho(a_{i})$ and $B_{i}=\rho(b_{i})$, as well as $C_{j}=\rho(c_{j})$. Proposition 2.5 implies that all the $A_{i}$ and $B_{i}$ are contained in the same orthogonal subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, which we can assume to be $\operatorname{PSO}(2)$ up to conjugating $\rho$. In particular, $[A_{i},B_{i}]=1$ for every index $i$.
Let’s first assume that there exists an index $i\in\{1,\ldots,g\}$ such that $A_{i}\neq 1$ or $B_{i}\neq 1$. We’ll consider the case where $A_{i}\neq 1$ (similar arguments apply when $B_{i}\neq 1$). Using the simple closed curves (2.1) from Lemma 2.6 and since $\rho$ is totally elliptic, we conclude that $A_{i}^{-1}[A_{i},B_{i}]C_{j}^{-1}A_{i}C_{j}$ is elliptic for every index $j$. Because $[A_{i},B_{i}]=1$, we even have that $[A_{i}^{-1},C_{j}^{-1}]$ is elliptic. As both $A_{i}$ and $C_{j}$ are elliptic by total ellipticity of $\rho$, Lemma 2.1 implies that $[A_{i},C_{j}]=1$ and thus $C_{j}\in\operatorname{PSO}(2)$ because $A_{i}\neq 1$. This holds for every index $j$ and therefore the image of $\rho$ is contained in $\operatorname{PSO}(2)$.
Now, let’s consider the case where $A_{i}=B_{i}=1$ for every $i\in\{1,\ldots,g\}$. Since we’re assuming $\rho$ to be reduced, we must have $n\geq 2$. Using the simple closed curve (2.2) from Lemma 2.6, we deduce that $B_{i}C_{k}B_{i}^{-1}C_{j}B_{i}A_{i}^{-1}B_{i}^{-1}C_{k}^{-1}B_{i}^{-1}C_{j}^{-1}$ is elliptic for every index $i$ and $j>k$. Since $A_{i}=B_{i}=1$, we even have $[C_{k},C_{j}]=1$ for every $j>k$ in $\{1,\ldots,n\}$. This implies that $\rho$ is orthogonal for the same reasons as before. ∎
Propositions 2.5 and 2.8 say that the only examples of reduced totally elliptic representations $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ when $\Sigma$ has positive genus and an arbitrary number of punctures are orthogonal representations.

### v2

@@ 1.5. Acknowledgements
2. Totally elliptic representations in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$
@@ 2.1. Overview
This section contains the proof of Theorem A which unfolds over the next subsections. We start by showing that when the genus of $\Sigma$ is at least $1$, then the only totally elliptic representations are orthogonal representations (Section 2.2). The arguments mostly rely on the observation that a commutator of elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is itself elliptic only when it is trivial (Lemma 2.1). We then briefly recall what DT representations are (Section 2.3) and prove that a totally elliptic genus-0 surface group representation that is not orthogonal must be a DT representation. We start with the case of spheres with three or four punctures (Sections 2.3.1 and 2.3.2), and then proceed by induction to prove the statement for spheres with an arbitrary number of punctures (Section 2.3.3).
@@ 2.2. Positive genus
We start by proving that a totally elliptic representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ of a surface $\Sigma$ of genus $g\geq 1$ is necessarily an orthogonal representation (Propositions 2.5 and 2.8). Most of the reasoning will rely on the following standard results about $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, which we include for the sake of completeness.
According to Definition 1.1, the elliptic elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ fall into two categories: the regular elliptic elements and the identity. The regular elliptic elements are those whose trace is smaller than two in absolute value. Geometrically, they act on the hyperbolic plane $\operatorname{\mathbb{H}}$ by rotations around their unique fixed point. The subset of regular elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is homeomorphic to the open ball $(0,2\pi)\times\operatorname{\mathbb{H}}$, where the first factor is the rotation angle and the second is the fixed point (the centre of rotation). A regular elliptic element $A\in\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is contained in a unique maximal compact subgroup which is a conjugate of $\mathrm{PSO}(2)$. It is therefore abelian and consists of all the elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ which fix the unique fixed point of $A$. It also coincides with the centralizer of $A$ in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
Lemma 2.1 (see e.g. [Bea83, Theorem 7.39.2]).
If $A\in\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is regular elliptic and $X\in\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is any element, then the commutator
@@
is either trivial or hyperbolic. Furthermore, the commutator is trivial if and only if $X$ is elliptic and fixes the unique fixed point of $A$.
@@
First, if $[A,X]=1$, then $X$ belongs to the centralizer of $A$, which means that $X$ is elliptic and fixes the unique fixed point of $A$. Now, if $[A,X]\neq 1$ and $o$ denotes the unique fixed point of $A$, then $XA^{-1}X^{-1}$ is regular elliptic with fixed point $X.o\neq o$. Note that if $\alpha$ is the rotation angle of $A$, then the rotation angle of $XA^{-1}X^{-1}$ is $2\pi-\alpha$. Let $\ell$ denote the unique geodesic line in $\operatorname{\mathbb{H}}$ through $o$ and $X.o$. Similarly, let $\ell_{1}$ and $\ell_{2}$ denote the geodesic lines obtained by rotating $\ell$ anti-clockwise by an angle $\alpha/2$ around $o$, respectively $X.o$.
@@
Lemma 2.1 plays a central role in our discussion because it tells us that the commutator of two elliptic elements is elliptic if and only if it is trivial. It turns out that two simple closed curves with geometric intersection number one can be lifted to specific fundamental group elements whose commutator represents a simple closed curve too. We must however be careful, because a simple closed curve on $\Sigma$ is represented by many elements in $\pi_{1}\Sigma$, all conjugate to each other.
@@
Let $\alpha$ and $\beta$ be two simple closed loops on $\Sigma$ intersecting once at $x\in\Sigma$ and such that the geometric intersection number of their free homotopy classes is one. We denote by $a$ and $b$ the fundamental group elements with basepoint $x$ associated to $\alpha$ and $\beta$. If $[a,b]$ is a non-trivial fundamental group element (which is always the case if $\Sigma$ is not a closed torus), then it represents a simple closed curve on $\Sigma$.
@@
Enlarge the two loops $\alpha$ and $\beta$ by replacing them with a thin ribbon. The union of the two ribbons is homeomorphic to a torus with one boundary component which is an embedded circle in $\Sigma$. The free homotopy class of that circle is non-trivial as long as $\Sigma$ is not a closed torus, and it corresponds to the free homotopy class determined by the fundamental group element $[a,b]$. ∎
As we are about to see, Lemmas 2.1 and 2.2 together provide the main obstruction for constructing totally elliptic surface group representations. Let us start by introducing some notation. We pick a geometric generating family for $\pi_{1}\Sigma$ which consists in $2g+n$ generators $a_{1},\ldots,a_{g}$, $b_{1},\ldots,b_{g}$, and $c_{1},\ldots,c_{n}$ satisfying the unique relation
@@
A possible choice of geometric generators is given by the homotopy classes of the loops illustrated on Figure 1.
[…]
@@
We adopt the convention of concatenating loops by starting with the rightmost loop. More precisely, if $\gamma_{1}$ and $\gamma_{2}$ are two loops on $\Sigma$ based at the same point, then their concatenation $\gamma_{1}\gamma_{2}$ is the loop obtained by going along $\gamma_{2}$ first, and then along $\gamma_{1}$. This rule applies for fundamental group multiplications too.
@@
If $\Sigma$ is a surface of genus $g\geq 2$ and $i\neq j$ are two distinct indices in $\{1,\ldots,g\}$, then the following fundamental group elements represent simple closed curves on $\Sigma$:
@@
Each pair of fundamental group elements appearing in one of the commutators above represents two simple closed curves on $\Sigma$ with geometric intersection number one. By Lemma 2.2, we can thus expect their commutators to represent simple closed curves too. This is however not automatic because of the issues we mentioned before stating Lemma 2.2. Thankfully, the way we picked the generators $a_{1},\ldots,a_{g}$ and $b_{1},\ldots,b_{g}$ ensures that each commutator does indeed represent a simple closed curve. This can be seen by drawing the corresponding loops using the generators of Figure 1 and the concatenation conventions described in Remark 2.3, like in the figure below.
[Uncaptioned $i$$j$$[b_{i}^{-1}b_{j}^{-1},a_{1}]$
@@
Assume that $\Sigma$ is a closed surface of genus $g\geq 1$. If $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is a totally elliptic representation, then $\rho$ is an orthogonal representation.
@@
We shall write $A_{i}=\rho(a_{i})$ and $B_{i}=\rho(b_{i})$. If $g=1$, then the image of $\rho$ is generated by $A_{1}$ and $B_{1}$ which satisfy $[A_{1},B_{1}]=1$. So, the image of $\rho$ is generated by two commuting elliptic elements, hence contained in a compact subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
Let us now consider the case $g\geq 2$. If $\rho$ is the trivial representation, then it is orthogonal. If not, then there exists an index $i\in\{1,\ldots,g\}$ such that either $A_{i}\neq 1$ or $B_{i}\neq 1$. We will look at the case where $A_{i}\neq 1$ (the case where $B_{i}\neq 1$ is identical). Let $j\neq i$ be another index. Since $\rho$ is totally elliptic, it maps all the simple closed curves from Lemma 2.4 to the identity because of Lemma 2.1. In particular, we have $[A_{i},B_{i}]=1$ and $[A_{i},B_{i}^{-1}B_{j}^{-1}]=1$, implying that $[A_{i},B_{j}]=1$. Since $[A_{i},B_{i}]=1$, up to conjugating $\rho$, we may assume that $A_{i}$ and $B_{i}$ belong to $\operatorname{PSO}(2)$. From $[A_{i},B_{j}]=1$ and our assumption that $A_{i}\neq 1$, we conclude that $B_{j}\in\operatorname{PSO}(2)$.
We also infer from Lemmas 2.1 and 2.4 that $[A_{j},B_{j}]=[A_{j},B_{j}^{-1}B_{i}^{-1}]=[A_{j},B_{i}A_{i}^{-1}B_{j}]=1$. If $B_{j}\neq 1$, then we immediately conclude that $A_{j}\in\operatorname{PSO}(2)$ from $[A_{j},B_{j}]=1$. Since $[A_{j},B_{j}]=1$ and $[A_{j},B_{j}^{-1}B_{i}^{-1}]=1$, we deduce that $[A_{j},B_{i}]=1$ which also implies $A_{j}\in\operatorname{PSO}(2)$ as soon as $B_{i}\neq 1$. If both $B_{i}=1$ and $B_{j}=1$, then we alternatively observe that $[A_{i}^{-1},A_{j}]=[B_{i}A_{i}^{-1}B_{j},A_{j}]=1$ which gives $A_{j}\in\operatorname{PSO}(2)$ as well, because $A_{i}\neq 1$.
@@
The first fundamental group element in (2.1) can be rewritten as $b_{i}a_{i}^{-1}b_{i}^{-1}\cdot c_{j}^{-1}a_{i}c_{j}$, which is the product of a conjugate of $a_{i}^{-1}$ and a conjugate of $a_{i}$. A similar observation holds for the second fundamental group element in (2.1) which is equal to $b_{i}a_{i}b_{i}a_{i}^{-1}b_{i}^{-1}\cdot c_{j}^{-1}b_{i}^{-1}c_{j}$ and is thus a product of conjugates of $b_{i}$ and $b_{i}^{-1}$. We can draw closed loops that represent these two fundamental group elements, as well as the one in (2.2), using the generators of Figure 1 and the concatenation rules of Remark 2.3, like in the figure below. This shows that they all three represent simple closed curves on $\Sigma$.
[Uncaptioned $i$$j$$a_{i}^{-1}[a_{i},b_{i}]c_{j}^{-1}a_{i}c_{j}$[Uncaptioned $i$$j$$k$$b_{i}c_{k}b_{i}^{-1}c_{j}b_{i}a_{i}^{-1}b_{i}^{-1}c_{k}^{-1}b_{i}^{-1}c_{j}^{-1}$
@@
A surface group representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ is said to be reduced if it does not map any peripheral curve on $\Sigma$ to the identity. Equivalently and in terms of the geometric generators from Figure 1, $\rho$ is reduced if it does not map any of the generators $c_{1},\ldots,c_{n}$ to the identity.
Every non-reduced surface group representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ factorizes through a reduced representation $\rho^{\prime}\colon\pi_{1}\Sigma^{\prime}\to\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$, where $\Sigma^{\prime}$ is obtained by filling in the punctures of $\Sigma$ whose corresponding peripheral curves are mapped to the identity by $\rho$. More precisely, the inclusion $\Sigma\subset\Sigma^{\prime}$ induces a surjective morphism $p\colon\pi_{1}\Sigma\twoheadrightarrow\pi_{1}\Sigma^{\prime}$ and $\rho^{\prime}\circ p=\rho$. Note that if $\rho$ is totally elliptic, then so is $\rho^{\prime}$.
@@
Assume that $\Sigma$ is a surface of genus $g\geq 1$ and with $n\geq 1$ punctures. If $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is a reduced totally elliptic representation, then $\rho$ is orthogonal.
@@
As before, we shall write $A_{i}=\rho(a_{i})$ and $B_{i}=\rho(b_{i})$, as well as $C_{j}=\rho(c_{j})$. Proposition 2.5 implies that all the $A_{i}$ and $B_{i}$ are contained in the same orthogonal subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, which we can assume to be $\operatorname{PSO}(2)$ up to conjugating $\rho$. In particular, $[A_{i},B_{i}]=1$ for every index $i$.
Let us first assume that there exists an index $i\in\{1,\ldots,g\}$ such that $A_{i}\neq 1$ or $B_{i}\neq 1$. We will consider the case where $A_{i}\neq 1$ (similar arguments apply when $B_{i}\neq 1$). Using the simple closed curves (2.1) from Lemma 2.6 and since $\rho$ is totally elliptic, we conclude that $A_{i}^{-1}[A_{i},B_{i}]C_{j}^{-1}A_{i}C_{j}$ is elliptic for every index $j\in\{1,\ldots,n\}$. Because $[A_{i},B_{i}]=1$, we even have that $[A_{i}^{-1},C_{j}^{-1}]$ is elliptic. As both $A_{i}$ and $C_{j}$ are elliptic by total ellipticity of $\rho$, Lemma 2.1 implies that $[A_{i},C_{j}]=1$ and thus $C_{j}\in\operatorname{PSO}(2)$ because $A_{i}\neq 1$. This holds for every index $j$ and therefore the image of $\rho$ is contained in $\operatorname{PSO}(2)$.
Now, let us consider the case where $A_{i}=B_{i}=1$ for every $i\in\{1,\ldots,g\}$. Since we are assuming $\rho$ to be reduced, we must have $n\geq 2$. Using the simple closed curve (2.2) from Lemma 2.6, we deduce that $B_{i}C_{k}B_{i}^{-1}C_{j}B_{i}A_{i}^{-1}B_{i}^{-1}C_{k}^{-1}B_{i}^{-1}C_{j}^{-1}$ is elliptic for every index $i\in\{1,\ldots,g\}$ and every $j>k$ in $\{1,\ldots,n\}$. Since $A_{i}=B_{i}=1$, we obtain $[C_{k},C_{j}]=1$ for every $j>k$ in $\{1,\ldots,n\}$. This implies that $\rho$ is orthogonal for the same reasons as before. ∎
Propositions 2.5 and 2.8 say that when $\Sigma$ has positive genus and an arbitrary number of punctures, then every reduced totally elliptic representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is orthogonal. In order to finish the proof of Theorem A, we shall now prove that if the genus of $\Sigma$ is zero, then every reduced totally elliptic surface group representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is either orthogonal or a DT representation.